# Bayesian Model Parameter Learning in Linear Inverse Problems: Application in EEG Focal Source Imaging

### v1

@@ 1.1 Model Uncertainties in Inverse Problems
In applications like computerized tomography and magnetic resonance imaging [10, 11], well-established methods are employed to tackle unknown model parameters. In particular, for the estimation of signals of interest and model parameters, the corresponding bilinear optimization problems can be approached by convexifying the problem using methods such as linearization (lifting) [12, 13, 11] or by alternating minimization algorithms [14, 15]. However, optimizing both the primary unknowns and the model parameters can be time consuming, convergence is not always guaranteed, or several reinitializations may be required, and approximations for parameters that are deep in the model may not always work. Techniques that handle model uncertainties with the help of machine learning (ML) and data-driven approaches [16, 17, 18] represent a new direction to accelerate computations and improve estimates in inverse problems [19, 20, 21]. However, unlike in many image and pattern recognition problems [22, 23], in inverse problems there is often a limited or no access to ground truth data to train ML algorithms and deep neural networks [10, 23]. In this study, we revisit the application of machine learning to severely ill-posed inverse problems when no access to ground truth is possible. Specifically, we generate training data stemming from physics-based models and integrate this domain-specific knowledge into the inverse problem formulation and recover both the primary unknown and extract an estimate for the (unknown) model parameter. We approach the problem from a Bayesian perspective, which facilitates the design of optimization algorithms under model uncertainty [24, 25].
As an application, we will focus on imaging brain sources from EEG recordings, due to its importance as a low-cost diagnostic tool and the inherent mathematical challenges and associated model uncertainties. In particular, we study imaging under the influence of erroneous skull conductivity (too high or too low), which has been shown to significantly affect the imaging solution [26, 27, 28, 29, 30, 31, 32]. Subsequently, the solved conductivity value could be used in further EEG studies and in the optimization/planning of transcranial electrical brain stimulation treatment [33, 34, 35]. We remark that the developments of this paper are not limited to EEG source imaging but can also be used, for example, in other imaging and deconvolution problems.
@@ 1.2 Bayesian Approximation Error Approach and Our Contributions
In this work, we aim to bring Bayesian statistics, which is not often explicitly used or contributed, into machine learning and provide methodological ways to efficiently solve linear inverse problems with model parameter uncertainties. In particular, our developments are based on the well-known Bayesian approximation error (BAE) approach [24, 36] in which a standard (approximate) observation model is used in the imaging accompanied with an additive error term that takes into account the effects of the approximation. Similarly as in [37, 36], we first derive a posterior distribution which simultaneously predicts the primary unknown signal and a low-order representation of the model-induced error. For the representation, we employ the top-eigenvectors of the modeling error covariance, obtained through Monte Carlo simulations. In addition to this, we derive an analytical expression for the modeling error covariance to investigate the connections between the error, model parameter, and primary signal. The provided insight allowed us to propose other Bayesian strategies to infer the model parameter.
In the application domain, we evolve our previous work, in which we demonstrated that BAE can improve source localization when skull conductivity was unknown [31], and our preliminary study on simultaneous estimation of focal source activity and skull conductivity from EEG recordings [38]. Here, with the help of simulated EEG data that corresponded to focal source activity in the somatosensory brain area, we demonstrate the potential of our approach to reconstruct the underlying focal sources and low-order estimates for the errors induced by unknown skull conductivity. Subsequently, the estimated errors are used to approximate the skull conductivity.
@@ 2.1 Bayesian Inference in Linear Inverse Problems
Let us assume the discrete observation model
@@
where $v\in\mathbb{R}^{m}$ are the measurements, $m$ is the number of measurements, $A(\sigma)\in\mathbb{R}^{m\times n}$ ($m\ll n$) is the system matrix (discrete forward model) that depends on a model parameter $\sigma\in\mathbb{R}^{N_{\sigma}}$, $x\in\mathbb{R}^{n}$ is the distributed signal (e.g. amplitudes in $n$ locations) that we wish to solve, and $e$ is the measurement noise, which is modeled as Gaussian $e\sim\mathcal{N}(e_{*},\Gamma_{e})$ with mean $e_{*}$ and covariance $\Gamma_{e}$. The model parameter $\sigma$ is usually (unrealistically) considered known. However, in this paper, we treat $\sigma$ as unknown, and subsequently estimate it alongside our primary unknown $x$.
In the Bayesian framework, the inverse solution is the posterior density
@@
For the model (1), the likelihood can be written as
@@ 2.2 Bayesian Machine Learning
Machine learning in inverse problems relies on training a model with known data, $D=\{(x_{s},y_{s})\}_{s=1:S}$, in order to predict underlying signals $x$ given a new set of observations, $y$. In Bayesian ML, model parameter $\sigma$ is a random variable with distribution $\pi(\sigma|D)$ and the likelihood is expressed as
@@ 3.1 Overview
To ease the model parameter estimation, we split the problem in two parts. First, we rewrite the observation model (1) with the help of an approximate model $A_{0}=A(\sigma_{0})$, where $\sigma_{0}$ has fixed values, in such a way that we remove the non-linearity with respect to the model parameter as
$v=A_{0}x+\varepsilon+e.$ (5)
Here, we added an error term (to compensate for the simplification),
$\varepsilon(x,\sigma)=A(\sigma)x-A_{0}x,$ (6)
which is referred to as approximation error, $\varepsilon\in\mathbb{R}^{m}$ [24, 36]. This linear error term enables the simplification of the likelihood function (linearization of the problem) under the assumption that this error term can be expressed as a linear combination of few basis functions, i.e. $\varepsilon=W\alpha+\varepsilon_{*}$, where $\varepsilon_{*}$ is the mean of $\varepsilon$, and $W\in\mathbb{R}^{m\times p}$ ($p<m$) are basis functions (obtained as in [37, 36]), we can infer $x$ and $\alpha$ based on Bayes’ rule, $\pi(x,\alpha|v)\propto\pi(v|x,\alpha)\pi(x,\alpha)$.
@@
Then, one way to obtain parameter estimates is by maximizing the conditional probability distribution $\hat{\sigma}:=\max_{\sigma}\pi(\sigma|\alpha)$, considering a joint Gaussian for $(\sigma,\alpha)$ as has been done previously [37]. However, this approach has limitations, since it considers that $\sigma$ and $\alpha$ are sensibly related through a linear relationship (or at least monotonic relationship), which may not be valid, and the approach also ignores the fact that $\alpha$ may depend on the signal $x$. This approximation can work usually well when the difference $\sigma-\sigma_{0}$ is small.
@@
In a more general case, we should consider that there is a relationship $f(\alpha,x)=\sigma$. Since the analytical expression of the function $f$ is often unknown (or too complicated to be derived), we can use non-parametric methods, such as a Gaussian process, for the Bayesian inference of the model parameter. Overall, as we will see later by approximating (6) and obtaining statistical knowledge through training data, we can recover meaningful estimates $\hat{\sigma}$ under certain conditions.
3.2 Approximation Error through Basis Functions
A feasible set of basis functions that assigns most of the variance of $\varepsilon$ in the first few terms can be chosen based on the eigenvalue decomposition of the approximation error covariance $\Gamma_{\varepsilon}=\mathbb{E}[(\varepsilon-\varepsilon_{*})(\varepsilon-\varepsilon_{*})^{\mathrm{T}}]$, where $\varepsilon_{*}$ is the mean of the modeling error. In particular, $\Gamma_{\varepsilon}$ can be decomposed according to
$\Gamma_{\varepsilon}=\Sigma_{k=1}^{m}\lambda_{k}w_{k}w_{k}^{\mathrm{T}}=W\Lambda W^{\mathrm{T}},$ (7)
where $\lambda_{k}\in\mathbb{R}^{m}$ are the eigenvalues and $w_{k}\in\mathbb{R}^{m}$ are the eigenvectors of $\Gamma_{\varepsilon}$ [37, 36]. Based on this decomposition, we have that
$\varepsilon-\varepsilon_{*}\;\in\;\mathrm{span}\{w_{1},\ldots,w_{k}\}.$ (8)
Hence, the approximation error can be expressed as
where $\varepsilon^{\prime}=\Sigma_{k=1}^{p}\alpha_{k}w_{k}$, $\varepsilon^{\prime\prime}=\Sigma_{j=p+1}^{m}\beta_{j}w_{j}$, with $\mathbb{E}[\varepsilon^{\prime}]=0$, $\mathbb{E}[\varepsilon^{\prime\prime}]=0$ and the coefficients $\alpha_{k}$ and $\beta_{j}$ are given by the inner products $\alpha_{k}=<\varepsilon-\varepsilon_{*},w_{k}>$ and $\beta_{j}=<\varepsilon-\varepsilon_{*},w_{j}>$ [37, 36]. The observation model (5) can be rewritten as
$v=A_{0}x+\varepsilon_{*}+W_{p}\alpha+\varepsilon^{\prime\prime}+e,$ (10)
where $\varepsilon^{\prime}=W_{p}\alpha$, $W_{p}=[w_{1},w_{2},...,w_{p}]\in\mathbb{R}^{m\times p}$ contains the first $p$ eigenvectors and $\alpha=(\alpha_{1},\alpha_{2},...,\alpha_{p})^{\mathrm{T}}\in\mathbb{R}^{p}$ [37, 36]. Note that $\varepsilon^{\prime\prime}=Q\beta=QQ^{\mathrm{T}}\varepsilon$ where $Q=[w_{p+1},\ldots,w_{m}]\in\mathbb{R}^{(m-p)\times m}$ and $\mathrm{var}(\alpha_{p})=\lambda_{p}$. In practice, the coefficients of $W$ can be learned from training data obtained from model distributions $\pi(x)$ and $\pi(\sigma)$ or semi-analytically as we will explain next.
3.2.1 Semi-Analytical Covariance Matrix for Approximation Error
The approximation error (6) depends on both, $\sigma\in\mathbb{R}^{+}$ and signal $x$. Therefore, to understand how the linear and non-linear effects of $\sigma$ and $x$ are incorporated in the basis functions $W$, we express the error covariance semi-analytically with the help of the Taylor expansion of the error (6). In particular, the Taylor expansion around the mean $\sigma_{*}$ is
$\varepsilon(x,\sigma)=A(\sigma_{*})x-A_{0}x+\sum_{k}\frac{\partial^{k}A(\sigma_{*})}{\partial\sigma^{k}}(\sigma-\sigma_{*})^{k}x=\sum_{k}J_{k}(\sigma_{*})z_{k}(\sigma)x,$ (11)
where $J_{k}=\frac{\partial^{k}A(\sigma)}{\partial\sigma^{k}}|_{\sigma=\sigma_{*}}\in\mathbb{R}^{m\times n}$, $z_{k}(\sigma)=(\sigma-\sigma_{*})^{k}\mathrm{1}_{n}$, $\mathrm{1}_{n}$ is a vector of ones of size $n$, and we have assumed $\sigma_{*}=\sigma_{0}$ for simplicity. Since $z_{k}$ and $x$ are statistically independent, according to [49], we have that the covariance matrix of a product of two independent variables is
$\mathrm{cov}[xz_{k}]=\mathrm{cov}[x]\mathrm{cov}[z_{k}]+\mathrm{cov}[x]\mathbb{E}[z_{k}]\mathbb{E}[z_{k}]^{\mathrm{T}}+\mathrm{cov}[z_{k}]\mathbb{E}[x]\mathbb{E}[x]^{\mathrm{T}}.$ (12)
Therefore, the error covariance can be expressed as
$\Gamma_{\varepsilon}=\sum_{k}J_{k}\mathrm{cov}[xz_{k}]J_{k}^{\mathrm{T}}.$ (13)
If we assume $x\sim\mathcal{N}(0,\gamma I_{n})$, then we have that
$\Gamma_{\varepsilon}=\gamma\sum_{k}J_{k}(\mathrm{cov}[z_{k}])+\mathbb{E}[z_{k}]\mathbb{E}[z_{k}]^{\mathrm{T}})J_{k}^{\mathrm{T}}.$ (14)
Based on the previous expression of the error covariance, we can observe that if an i.i.d Gaussian prior for the signal $x$ is assumed, then $x$ only has a scaling effect on the error covariance.
In general, we need to analyze the properties of each studied problem to determine whether the matrix $J_{k}$ implicitly encodes information about $x$ (potentially coming from the modeling, e.g., linearization of a nonlinear problem). Thus, the inference of the model parameter $\sigma$ from the estimated error $\varepsilon^{\prime}$ (as in [37]) is feasible if the basis functions $W$ primarily reflect the error induced by the unknown model parameter $\sigma$. However, this is not always the case, and in some cases also the signal $x$ must be considered when estimating $\sigma$.
@@ 3.3 Bayesian Inference of Primary Signal and Approximation Error
For the simultaneous estimation of $x$ and $\alpha$ (and $\varepsilon\approx W\alpha+\varepsilon_{*}$), we have to construct the posterior model $\pi(x,\alpha|v)$. To obtain a computationally efficient solution, we make the technical approximation that $(x,\alpha,e,\varepsilon^{\prime\prime})$ are mutually Gaussian and uncorrelated [^1] We note that even though in reality approximation errors usually depend on $x$, this approximation often leads to very similar inverse solutions [24, 25, 36]. [37, 36]. Then, we obtain the approximate likelihood
$\tilde{\pi}(v|x,\alpha)\propto\exp\Big{(}-\frac{1}{2}(v-A_{0}x-\varepsilon_{*}-W_{p}\alpha-e_{*})^{\mathrm{T}}$
$\qquad(\Gamma_{\varepsilon^{\prime\prime}}+\Gamma_{e})^{-1}(v-A_{0}x-\varepsilon_{*}-W_{p}\alpha-e_{*})\Big{)},$ (15)
where $\Gamma_{\varepsilon^{\prime\prime}}=\Sigma_{p+1}^{m}\lambda_{j}w_{j}w_{j}^{\mathrm{T}}$. Thus, the posterior density becomes
$\tilde{\pi}(x,\alpha|v)\propto\mathrm{exp}\left(-\frac{1}{2}L_{\varepsilon^{\prime\prime}+e}\|v-A_{0}x-W\alpha-\varepsilon_{*}-e_{*}\|_{2}^{2}\right)\;\pi(x,\alpha),$ (16)
where the Cholesky factors are $(\Gamma_{\varepsilon^{\prime\prime}}+\Gamma_{e})^{-1}=L_{\varepsilon^{\prime\prime}+e}^{\mathrm{T}}L_{\varepsilon^{\prime\prime}+e}$. Then, we can estimate the maximum a posteriori (MAP) estimate
$(\hat{x},\hat{\alpha})\leftarrow\min_{x,\alpha}\left\{\|\frac{1}{2}L_{\varepsilon^{\prime\prime}+e}(A_{0}x+W_{p}\alpha+e_{*}+\varepsilon_{*}-v)\|_{2}^{2}-\ln(\pi(x,\alpha)\right\}.$ (17)

### v2

@@ 1.1 Model Uncertainties in Inverse Problems
In applications like computerized tomography and magnetic resonance imaging [10, 11], well-established methods are employed to tackle unknown model parameters. In particular, for the estimation of signals of interest and model parameters, the corresponding bilinear optimization problems can be approached by convexifying the problem using methods such as linearization (lifting) [12, 13, 11] or by alternating minimization algorithms [14, 15]. However, optimizing both the primary unknowns and the model parameters can be time consuming, convergence is not always guaranteed or several reinitializations may be required, and approximations for parameters that are deep in the model may not always work. Techniques that handle model uncertainties with the help of machine learning (ML) and data-driven approaches [16, 17, 18] represent a new direction to accelerate computations and improve estimates in inverse problems [19, 20, 21]. However, unlike in many image and pattern recognition problems [22, 23], in inverse problems there is often a limited or no access to ground truth data to train ML algorithms and deep neural networks [10, 23]. In this study, we revisit the application of machine learning to severely ill-posed inverse problems when no access to ground truth is possible. Specifically, we generate training data stemming from physics-based models and integrate this domain-specific knowledge into the inverse problem formulation and recover both the primary unknown and extract an estimate for the (unknown) model parameter. We approach the problem from a Bayesian perspective, which facilitates the design of optimization algorithms under model uncertainty [24, 25].
As an application, we will focus on imaging brain sources from EEG recordings, due to its importance as a low-cost diagnostic tool and the inherent mathematical challenges and associated model uncertainties. In particular, we study imaging under the influence of erroneous skull conductivity (too high or too low), which has been shown to significantly affect the imaging solution [26, 27, 28, 29, 30, 31, 32]. In numerical experiments, we used simulated EEG recordings generated by single dipole sources in the somatosensory cortex of the brain, as similar data sets (EEG and MEG) have previously been found to be suitable for skull conductivity calibration [33, 34, 35, 36, 37, 38]. Subsequently, the inferred conductivity value could be used in further EEG studies and in the optimization/planning of transcranial electrical brain stimulation treatment [39, 40, 41]. We remark that the developments of this paper are not limited to EEG source imaging but can also be used, for example, in other imaging and deconvolution problems.
@@ 1.2 Bayesian Approximation Error Approach and Our Contributions
In this work, we build on the growing use of Bayesian statistics in machine learning by providing methodological approaches to efficiently solve linear inverse problems with model parameter uncertainties. In particular, our developments are based on the well-known Bayesian approximation error (BAE) approach [24, 42] in which a standard (approximate) observation model is used in the imaging accompanied with an additive error term that takes into account the effects of the approximation. Similarly as in [43, 42], we first derive a posterior distribution which simultaneously predicts the primary unknown signal and a low-order representation of the model-induced error. For the representation, we employ the top-eigenvectors of the modeling error covariance, obtained through Monte Carlo simulations. In addition to this, we derive an analytical expression for the modeling error covariance to investigate the connections between the error, model parameter, and primary signal. The provided insight allowed us to propose other Bayesian strategies to infer the model parameter.
In the application domain, we evolve our previous work, in which we demonstrated that BAE can improve source localization when skull conductivity is unknown [31], and our preliminary study on simultaneous estimation of focal source activity and skull conductivity from EEG recordings [44]. Here, with the help of simulated EEG data that corresponded to focal source activity in the somatosensory brain area, we demonstrate the potential of our approach to reconstruct the underlying focal sources and low-order estimates for the errors induced by unknown skull conductivity. Subsequently, these estimates are used to approximate the skull conductivity.
@@
The paper is organized as follows. We first provide mathematical preliminaries in Section 2. In Section 3, we then present a general overview of the Bayesian Approximation Error (BAE) approach and how it can be exploited for model parameter estimation. In Section 4, we describe how the BAE and the proposed model parameter estimation techniques can be applied to a model problem in which the source activity of the brain (primary unknown) and the electrical conductivity of the human skull (model parameter) are estimated from EEG recordings. Section 5 describes the implementation of the numerical simulations followed by the results and discussion. Section 6 consists of conclusions and future work. Finally, the appendices provide further details on the theory and analysis and a step-by-step pipeline on how to implement the BAE approach with model parameter estimation.
@@ 2.1 Bayesian Inference in Linear Inverse Problems
We consider the discrete observation model
@@
where $v\in\mathbb{R}^{m}$ denotes the measurements, $m$ is the number of measurements, $A(\sigma)\in\mathbb{R}^{m\times n}$ ($m\ll n$) is the system matrix (discrete forward model) that depends on a model parameter $\sigma\in\mathbb{R}^{N_{\sigma}}$, $x\in\mathbb{R}^{n}$ is the distributed signal (e.g. amplitudes in $n$ locations) that we aim to solve. Finally, $e$ denotes the measurement noise that is modeled as Gaussian $e\sim\mathcal{N}(e_{*},\Gamma_{e})$ with mean $e_{*}$ and covariance $\Gamma_{e}$. The model parameter $\sigma$ is usually (unrealistically) considered known. However, in this paper, we treat $\sigma$ as unknown, and subsequently estimate it alongside our primary unknown $x$.
In the Bayesian framework, the solution to the inverse problem is the posterior density
@@
In this case, the likelihood can be formulated as
@@ 2.2 Bayesian Machine Learning
Machine learning in inverse problems relies on training a model with known data, $D=\{(x_{s},v_{s})\}_{s=1:S}$, in order to predict underlying signals $x$ given a new set of observations, $v$. In Bayesian ML, model parameter $\sigma$ is a random variable with distribution $\pi(\sigma|D)$, $x$ and $\sigma$ are statistically independent and the likelihood is expressed as
@@ 3.1 Overview
To ease the model parameter estimation, we rewrite the observation model (1) with the help of an approximate model $A_{0}=A(\sigma_{0})\in\mathbb{R}^{m\times n}$, where $\sigma_{0}$ has fixed values, and remove the non-linearity with respect to the unknown model parameter $\sigma$ as
$v=A_{0}x+\varepsilon+e,$ (5)
where the additive error term $\varepsilon\in\mathbb{R}^{m}$, referred to as approximation error [24, 42], accounts for the simplification. In particular, the approximation error term is related to the unknown parameter as
$\varepsilon(x,\sigma)=A(\sigma)x-A_{0}x.$ (6)
Furthermore, when the variability of the error term $\varepsilon$ is relatively low, it can be expressed as a linear combination of few basis functions, for example, as in [43, 42] by using the eigenvalue decomposition of the error covariance, $\Gamma_{\varepsilon}=\Sigma_{k=1}^{m}\lambda_{k}w_{k}w_{k}^{\mathrm{T}}$, where $\lambda_{k}>0$ (in descending order) are the eigenvalues and $w_{k}\in\mathbb{R}^{m}$ the eigenvectors. Now, we can write the approximation error as
$\varepsilon(x,\sigma)=\varepsilon_{*}+\varepsilon^{\prime}+\varepsilon^{\prime\prime}=W^{p}\alpha+\varepsilon_{*}+\varepsilon^{\prime\prime},$ (7)
where $\varepsilon_{*}=\mathbb{E}[\varepsilon]$ is the statistical mean of $\varepsilon$ and $W^{p}=[w_{1},\ldots,w_{p}]\in\mathbb{R}^{m\times p}$ ($m>p$) are the $p$ top eigenvectors. The term $\varepsilon^{\prime}=W^{p}\alpha$ represents the dominant part of the variability in the error, while $\varepsilon^{\prime\prime}$ is a less informative term described by the remaining $m-p$ eigenvectors with mean $\mathbb{E}[\varepsilon^{\prime\prime}]=0$ and covariance $\Gamma_{\varepsilon^{\prime\prime}}=\Sigma_{j=p+1}^{m}\lambda_{j}w_{j}w_{j}^{\mathrm{T}}$. The coefficient vector $\alpha\in\mathbb{R}^{p}$ represents the low-dimensional error components with mean $\mathbb{E}[\alpha]=0$ and diagonal covariance $\mathrm{cov}[\alpha]=\mathrm{diag}\{\lambda_{1},\ldots,\lambda_{p}\}$. More details on this error decomposition are given in Appendix A1. Now, by substituting the decomposed error into the observation model (5) we get
$v=A_{0}x+\varepsilon_{*}+W^{p}\alpha+\varepsilon^{\prime\prime}+e.$ (8)
In practical applications, the basis $W^{p}$ and corresponding statistics can be learned from training data sampled from the joint model distributions $\pi(x)$ and $\pi(\sigma)$, or computed semi-analytically [43, 42]. Then, we can infer simultaneously $x$ and $\alpha$ based on Bayes’ rule, $\pi(x,\alpha|v)\propto\pi(v|x,\alpha)\pi(x,\alpha)$.
@@
Consequently, one way to obtain parameter estimates is by maximizing the conditional probability distribution $\hat{\sigma}:=\max_{\sigma}\pi(\sigma|\alpha)$, considering a joint Gaussian for $(\sigma,\alpha)$ as has been done previously [43]. However, this approach has limitations, since it considers that $\sigma$ and $\alpha$ are sensibly related through a linear relationship (or at least monotonic relationship), which may not be valid, and the approach also ignores the fact that $\alpha$ depends on the signal $x$. This approximation can work well when the difference $\sigma-\sigma_{0}$ is small.
@@
In a more general case, we can consider a mapping from $(\alpha,x)\rightarrow\sigma$. Since the analytical expression of this mapping is often unknown (or too complicated to be derived), we can use non-parametric methods, such as a Gaussian process (GP), for the inference of the model parameter $\sigma$. Since $x$ and $\sigma$ are usually independent (this depends on the underlying physics of the problem), the only way $x$ can affect $\sigma$ in this model is through its interaction with the variable $\alpha$; in our case, this kind of interaction occurs because $\alpha$ depends on $x$ through equations (6) and (7). Hence, GP learns a statistical relationship, not a causal one. GP is effectively learning the inverse function of $\sigma$ from the structure imposed by $A(\sigma)$. Overall, as we shall see later by approximating (6) and obtaining statistical knowledge through training data, we can recover meaningful estimates $\hat{\sigma}$ under certain conditions.
@@ 3.2 Bayesian Inference of Primary Signal and Approximation Error
To jointly estimate both $x$ and $\alpha$, we need the posterior distribution $\pi(x,\alpha\mid v)$. To reduce computational complexity, we adopt an approximation in which the variables $(x,\alpha,e,\varepsilon^{\prime\prime})$ are treated as mutually Gaussian and uncorrelated [^1] We note that even though in reality approximation errors usually depend on $x$, this technical approximation often results in very similar inverse solutions [24, 25, 42]. [43, 42]. The resulting approximate likelihood is
$\tilde{\pi}(v|x,\alpha)\propto\exp\Big{(}$ $-\frac{1}{2}(v-A_{0}x-\varepsilon_{*}-W^{p}\alpha-e_{*})^{\mathrm{T}}$
$(\Gamma_{\varepsilon^{\prime\prime}}+\Gamma_{e})^{-1}(v-A_{0}x-\varepsilon_{*}-W^{p}\alpha-e_{*})\Big{)}.$
The corresponding posterior density is
$\tilde{\pi}(x,\alpha|v)\propto\mathrm{exp}\left(-\frac{1}{2}\|L_{\varepsilon^{\prime\prime}+e}\left(v-A_{0}x-W^{p}\alpha-\varepsilon_{*}-e_{*}\right)\|_{2}^{2}\right)\;\pi(x,\alpha),$ (10)
where $L_{\varepsilon^{\prime\prime}+e}$ comes from the Cholesky factorization of $(\Gamma_{\varepsilon}+\Gamma_{e})^{-1}$. Finally, the maximum a posteriori (MAP) estimate is obtained by solving:
$(\hat{x},\hat{\alpha})\leftarrow\arg\min_{x,\alpha}\left\{\frac{1}{2}\left\|L_{\varepsilon^{\prime\prime}+e}\left(A_{0}x+W^{p}\alpha+\varepsilon_{*}+e_{*}-v\right)\right\|_{2}^{2}-\log\pi(x,\alpha)\right\}.$ (11)
Next, we can use the point estimates $(\hat{x},\hat{\alpha})$ to infer $\sigma$ with the help of learning techniques.